Circular No. 7989

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1650–500 C. Sánchez-Fernández, Research and Scientific Support Department,

European Space Agency; C. Zurita and J. Casares, Instituto de Astrofísica de Canarias; A. J. Castro-Tirado, Instituto de Astrofísica de Andalucía; I. Bond, Mt. John University Observatory; and S. Brandt and N. Lund, Danish Space Research Institute, report: “We obtained time-series spectroscopy of the blackhole candidate XTE J1650–500 on June 10 using the UT4 telescope (+ FORS2) at the European Southern Observatory, Paranal. Fifteen spectra (range 575–731 nm; resolution 0.076 nm FWHM) showed a broad, double-peaked Hα emission line (EW = 10 Å) from the accretion disk and weak absorption lines from the companion star. Radial-velocity curves of the secondary star were measured by cross-correlation with several G/K-type template stars, all giving consistent results. A four-parameter sinusoidal fit to the secondary radial-velocity data was performed, yielding the following parameters for the best fit: period, 0.212 ± 0.001 days; semiamplitude of the radial-velocity curve of the companion star, 309 ± 4 km/s; systemic radial velocity, 19 ± 3 km/s; T_{0} = HJD 2452436.543 ± 0.001 day. The reduced χ^{2} for the fit was 1.034. The resulting mass function for the compact object is then M_{1}sin^{3}i/(1+M_{2}/M_{1})^{2} = 0.64 ± 0.03 M_{⊙}. If the more massive compact object is indeed a blackhole (i.e., M_{1} > 3 M_{⊙}), as indicated by the x-ray properties (IAUC 7707, 7715), and assuming a typical mass ratio of 0.1 for blackhole binaries, the inclination of the system must then be lower than 40^{o}. The equivalent width of the Hα emission showed a ∼ 40-percent modulation with the 0.21-day orbital period. The weakness of the absorption lines from the companion suggests that residual emission from the accretion disk still dominates the optical spectrum, and hence XTE J1650–500 had not yet reached true quiescence at the time of our observations.”

COMET P/2002 S1 (SKIFF) Additional observations and the following elliptical orbital elements

appear on MPEC 2002-S61:

T = 2002 Mar. 26.5669 TT  ω = 35.^{o}6546
e = 0.437531  Ω = 347.4633 }2000.0
q = 2.305629 AU  i = 27.6887
a = 4.099120 AU n^{o} = 0.1187594 P = 8.299 years
2002 October 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green